Solve the system of the equations. I need to know how to work this problem out.
1. -2x-6y=-2
2x +3y=5
(this is all one problem)
2. 4x+y=5
2x-5y =19

Answer by checkley71(8403)

-2X-6Y=-2
2X+3Y=5 NOW ADD THESE 2 EQUATIONS
-----------------------------------
-3Y=3
Y=3/-3
Y=-1 ANSWER.
-2X-6*-1=-2
-2X+6=-2
-2X=-2-6
-2X=-8
X=-8/-2
X=4 ANSWER.
----------------------------------------
4X+Y=5 NOW MULTIPLY THE OTHER EQUATION BY -2
-4X+10Y=-38
-----------------------
11Y=-33
Y=-33/11
Y=-3 ANSWER.
4X-3=5
4X=5+3
4X=8
X=8/4
X=2 ANSWER.